# Acyclic Conjunctive Regular Path Queries are no Harder than Corresponding Conjunctive Queries

### v1

@@ Acknowledgements.
This work is in part supported by SNSF 200021-231956. Suciu’s work is partially supported by NSF-IIS 2314527, NSF-SHF 2312195, NSF-III 2507117, and a Microsoft professorship.
@@ Appendix B Missing details in Section 3
We start with some definitions. A (multi-)hypergraph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ consists of a finite set of vertices $\mathcal{V}$ and a (multi-)set of edges $\mathcal{E}$ such that each edge is a subset of $\mathcal{V}$ (of arbitrary size). Given a hypergraph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ and a vertex $v\in\mathcal{V}$, we define the sets $\mathcal{E}_{v}=\{e\in\mathcal{V}\mid v\in e\}$ and $\mathcal{V}_{v}=\bigcup_{e\in\mathcal{E}_{v}}e$. The hypergraph obtained by the elimination of $v$ in $\mathcal{G}$ is defined as $\textsf{eliminate}(\mathcal{G},v)=(\mathcal{V}^{\prime},\mathcal{E}^{\prime})$, where $\mathcal{V}^{\prime}=\mathcal{V}\setminus\{v\}$ and $\mathcal{E}^{\prime}=(\mathcal{E}\setminus\mathcal{E}_{v})\cup\{\mathcal{V}_{v}\setminus\{v\}\}$.
@@ Appendix C Comparison with Existing Work
Figure 3. The edge-labeled graph used in Example C.1:

### v2

@@ Acknowledgements.
Suciu’s work was partially supported by NSF IIS 2314527, NSF SHF 2312195, NSF III 2507117, and a Microsoft professorship. Olteanu’s work is partially supported by SNSF 200021-231956.
@@ Appendix B Missing details in Section 3
We start with some definitions. A (multi-)hypergraph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ consists of a finite set of vertices $\mathcal{V}$ and a (multi-)set of edges $\mathcal{E}$ such that each edge is a subset of $\mathcal{V}$ (of arbitrary size). Given a hypergraph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ and a vertex $v\in\mathcal{V}$, we define the sets $\mathcal{E}_{v}=\{e\in\mathcal{E}\mid v\in e\}$ and $\mathcal{V}_{v}=\bigcup_{e\in\mathcal{E}_{v}}e$. The hypergraph obtained by the elimination of $v$ in $\mathcal{G}$ is defined as $\textsf{eliminate}(\mathcal{G},v)=(\mathcal{V}^{\prime},\mathcal{E}^{\prime})$, where $\mathcal{V}^{\prime}=\mathcal{V}\setminus\{v\}$ and $\mathcal{E}^{\prime}=(\mathcal{E}\setminus\mathcal{E}_{v})\cup\{\mathcal{V}_{v}\setminus\{v\}\}$.
@@ Appendix C Comparison with Existing Work
Figure 3. The edge-labeled graph used in Example C.1
[…]